OpenStudy (anonymous):

What is the first step in solving 4(x + 2) = 10 + x?

OpenStudy (anonymous):

4x+8=10+x 4x-x=10-8 3x=2 x=2/3

OpenStudy (anonymous):

@kristinlovesyou

OpenStudy (anonymous):

Thank you!